OpenStudy (anonymous):

Solve for x: 4x - 2 > 6x + 8. x > -5 x < -5 x > 5 x < 5

OpenStudy (anonymous):

@Michele_Laino , @triciaal , @texaschic101 , @Preetha

OpenStudy (firekat97):

firstly, you want to start off by collecting like terms.. so try doing so and post up what you get

OpenStudy (anonymous):

x<-5

OpenStudy (michele_laino):

more precisely we can subtract from both sides 4x, so we get: \[4x - 2 - 4x > 6x + 8 - 4x\] please simplify

OpenStudy (anonymous):

I want to see how it's don't, not just the answer

OpenStudy (texaschic101):

please do not give direct answers .....if you do, at least tell how you got it

OpenStudy (anonymous):

How would I simplify that?

OpenStudy (michele_laino):

it is simple, you have to sum similar terms, so we get this: \[\begin{gathered} 4x - 2 - 4x = - 2 \hfill \\ 6x + 8 - 4x = 2x + 8 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I don't think I did it right.

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Wow, your awesome, thanks so much!

OpenStudy (michele_laino):

:)